General Instruction: Attempt all of the following questions 

1. Explain why economics deals with efficient utilization of scarce resources. 

2. Define scarcity, tradeoffs and opportunity cost. Give three examples of important tradeoffs that you face in your life 

3. Imagine a society that produces military goods and consumer goods, which we will call “guns” and “butter.”

a. Draw a production possibilities frontier for guns and butter. Explain why it most likely has a bowed out shape.

b. Show a point that is impossible for the economy to achieve. Show a point that is feasible but inefficient.

4. Assume a hypothetical consumer consumes good X and good Y. The price of good X is 1 and price of good Y is 3 and the consumer budget is birr 10 for the two goods. Where: QX is quantity of good X, QY is quantity of good Y and TUX and TUY is total utility from consuming good X and good Y respectively. 

Q TUX MUx MUx/Px TUY MUy MUy/Py

0 0 0

1 10 24

2 19 45

3 27 63

4 34 78

5 40 87

6 44 90

      Based on the given information, answer the following questions.

A. Compute the marginal utility of the two goods

B. At what amounts of consumption does diminishing marginal utility starts to occur for the two goods?

C. Determine the quantities of the two goods that the consumer should buy in order to maximize his total utility.

6. Explain the law of variable proportion.

7. Explain the reason why, in the short run, the fixed cost is said to be unavoidable cost 

8. As a firm procures more out puts the average variable cost become closer to the average total cost of production, explain this concept 

9. Suppose that a competitive firm’s total cost of producing output q is given by 

TC= 3+3Q+q2. Assume that the market price for the firm’s product is $9.

a. What level of output will the firm produce?

b. Will the firm be earning positive, negative or zero profit? 

10. A farmer grows a bushel of wheat and sells it to a miller for $1.00. The miller turns the wheat into flour and then sells the flour to a baker for $3.00. The baker uses the flour to make bread and sells the bread to an engineer for $6.00. The engineer eats the bread. What is the value added by each person? What is GDP?

11. Consider an economy that produces and consumes bread and automobiles. In the following table are data for two different years.

      Year                       Year

      2000                       2010

 

Automobiles    100         $50,000    120       $60,000

Bread       500,000       $10      400,000     $20

Using the year 2000 as the base year, compute the following statistics for each year: 

 

a) nominal GDP, 

b) real GDP, 

c) GDP deflator 

d) CPI.



1
Expert's answer
2021-02-17T14:52:01-0500
  1. Because economics is the science of efficient use of resources, which constantly gives answers to the following questions: what, how and for whom to produce? The answers to these questions are possible only through the effective use of limited resources.
  2. Compromise denotes an option that we give up in order to get what we want. The opportunity cost is the cost of the second best alternative that needs to be made in order to make a choice. Deficit — shortage; insufficiency of something. Examples of compromise: reduce costs for future investments, save for future expenses, choose the mode of transport for shipping products.
  3. a.



The shape of the production capacity curve shows the price of one product expressed in an alternative quantity of another.

b. point M is a point that is impossible for the economy to achieve. The point M is unreachable with the available amount of resources and this technology.

An inefficient point can be reached if resources are used inefficiently.

4.

A. MUX/MUY = 3

QX + 3QY =10

MUX=3

MUY=1

B. Qy=2, Qx=4

C. "\frac{MUx}{Px}=\frac{MUy}{Py}"

"\frac{MUX}{1}=\frac{MUy}{3}"

MUX=1

MUY=3

MUX+MUY=1+3=4

6.

The law of decreasing productivity is also called "the law of variable proportions" or " the law of changing proportions». According to this law, the increase in production achieved with a uniform increase in the variable factor, after reaching a certain level, will decrease as the ratio between the variable and constant factors increases.

7.As a rule, labor is a variable factor, and capital is a constant factor in the short term. Short-term: fixed costs have already been paid and are not refundable.

8.As production increases, the average variable costs first fall (positive economies of scale), reach their minimum, and then, under the law of diminishing returns, begin to grow.

9.

a. MR=MC=P

(3 + 3Q + q2)'=9

3+2q=9

q=3

b. Pr=TR-TC=9*3-(3+3*3+32)=27-21=6 profit

10.

A farmer added $1. The miller added $2 (3-1).The baker added $3 (6-3)

Gross domestic product is the market value of all goods and services produced in a country for final use, that is, intended for consumption, and not for the production of other goods or services.

11.



"GDR n=\Sigma Pt\times Qt"

"GDR r=\Sigma P0\times Qt"

"GDR deflator=\frac{\Sigma(Q1\times P1}{\Sigma(Q1\times P0}"

"InflationRate=\\frac{(CurrentPeriodCPI\u2212PriorPeriodCPI\u200b)}{PriorPeriodCPI}"